# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 19 Oct 2023 v1

@@ 7 Conclusion
Recognizing the shortcomings of existing MU approaches, we introduce the innovative concept of weight saliency in MU, leading to the SalUn framework. This proposed saliency unlearning has showcased its effectiveness in addressing the limitations of current MU methods, and applying to both image classification and generation tasks. As a notable application, we show that SalUn stays effective in preventing stable diffusion from generating images with harmful content, even suffering inappropriate image prompts.
@@ B.1 Additional training and unlearning settings
For the Retrain method, training is conducted over 182 epochs using the SGD optimizer with a cosine-scheduled learning rate initialized at 0.1. For both FT and RL, training spans 10 epochs with a fixed learning rate of 0.1. GA’s training settings involve a 5-epoch learning rate search within the interval [1e-6, 1e-4]. In the case of IU, we explore the parameter $\alpha$ associated with the woodfisher Hessian Inverse approximation within the range [1, 20]. For $\ell_{1}$-sparse, a learning rate search for the parameter $\gamma$ is executed within the range [1e-6, 1e-4], while maintaining a constant learning rate of 0.1. For the BS method, the FGSM step size is defined as 0.1. Both BS and BE methods involve a 10-epoch learning rate search in the interval [1e-6, 1e-4]. Lastly, for SalUn and SalUn-soft, training is carried out over 10 epochs with a set learning rate of 0.1.
@@
For DDPM, the unlearning settings are as follows: For Retrain, it undergoes training for 80,000 iterations with Adam and a learning rate of 1e-4. The batch size is set to 128. In the case of SalUn, it is trained for 1,000 iterations with a learning rate of 1e-4, $\alpha$ set to 1e-3, and a batch size of 128. The sparsity of weight saliency is maintained at 50%. The sampling settings include 1,000 time steps and a conditional scale of 2.0.
For SD, the unlearning settings are as follows: For SalUn, it undergoes training for 5 epochs with Adam using a learning rate of 1e-5. The $\alpha$ value is set at 0.5, with a batch size of 8. The sparsity of weight saliency is again set at 50%. The sampling settings involve the use of DDIM, 100 time steps, and a conditional scale of 7.5.
@@ Appendix C Broader Impacts and Limitations
Figure A1: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red border.(More results will be shown in Fig. A2 and Fig. A3)
@@
Figure A2: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red border. (Extended results from Fig. A1, more results will be shown in Fig. A3)
@@
Figure A3: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red border. (Extended results from Fig. A2)

### 06 Nov 2023 v2

@@ 7 Conclusion
Recognizing the shortcomings of existing MU (machine unlearning) approaches, we introduce the innovative concept of weight saliency in MU, leading to the SalUn framework. The proposed saliency unlearning has showcased its effectiveness in addressing the limitations of current MU methods, and applying to both image classification and generation tasks. As a notable application, we show that SalUn stays effective in preventing stable diffusion from generating images with harmful content, even when given inappropriate image prompts. As other future research directions, we will explore the theoretical foundations of weight saliency-aided unlearning and consider other MU applications in generative AI, such as large language models.
@@ B.1 Additional training and unlearning settings
For the Retrain method, training is conducted over $182$ epochs using the SGD optimizer with a cosine-scheduled learning rate initialized at $0.1$. For both FT and RL, training spans 10 epochs with a fixed learning rate of $0.1$. GA’s training settings involve a 5-epoch learning rate search within the interval $10^{-6},10^{-4}]$. In the case of IU, we explore the parameter $\alpha$ associated with the woodfisher Hessian Inverse approximation within the range $[1,20]$. For $\ell_{1}$-sparse, a learning rate search for the parameter $\gamma$ is executed within the range $10^{-6},10^{-4}]$, while maintaining a constant learning rate of 0.1. For the BS method, the FGSM step size is defined as 0.1. Both BS and BE methods involve a 10-epoch learning rate search in the interval $10^{-6},10^{-4}]$. Lastly, for SalUn and SalUn-soft, training is carried out over 10 epochs with a set learning rate of 0.1.
@@
For DDPM, the unlearning settings are as follows: For Retrain, it undergoes training for 80,000 iterations with Adam and a learning rate of $10^{-4}$. The batch size is set to 128. In the case of SalUn, it is trained for 1,000 iterations with a learning rate of $10^{-4}$, $\alpha$ set to $10^{-3}$, and a batch size of 128. The sparsity of weight saliency is maintained at 50%. The sampling settings include 1,000 time steps and a conditional scale of 2.0.
For SD, the unlearning settings are as follows: For SalUn, it undergoes training for 5 epochs with Adam using a learning rate of $10^{-5}$. The $\alpha$ value is set at 0.5, with a batch size of 8. The sparsity of weight saliency is again set at 50%. The sampling settings involve the use of DDIM, 100 time steps, and a conditional scale of 7.5.
@@
Refer to caption Refer to caption Refer to caption (a) UA for classification (b) MIA for classification (c) TA for classification
Figure A1: Performance of SalUn with different saliency sparsity vs Retrain for classification. The settings follow Table 1. Points above Retrain indicate over-forgetting, while points below Retrain represent under-forgetting.
Ablation study on sparsity choice of weight saliency.
Recall from (3) that the choice of the weight saliency sparsity threshold could be a crucial hyperparameter in our approach. In our experiments, we have set the default sparsity threshold to 50%. Fig. A1 and Fig. A2 provide a more detailed examination of the sparsity ratio. Specifically, Fig. A1-(a), (b) and (c) present the performance of random data forgetting in image classification using ResNet-18 on CIFAR-10 (with the experiment setup same as Table 1). Fig. A2-(a) and (b) present the performance of class-wise forgetting for image generation using DDPM on CIFAR-10 with the same setting as Fig 4. In both cases, the performance of Retrain is provided for comparison.
Refer to caption Refer to caption (a) UA vs. sparsity (b) FID vs. saprsity
Figure A2: Performance of SalUn with different saliency sparsity vs Retrain in the tas of image generation. The settings follow Fig 4.
In the context of MU for image classification, choosing a 50% sparsity threshold is a reasonable option. A higher sparsity of the weight saliency may result in under-forgetting, as evidenced by the lower unlearning accuracy (Jia et al., 2023) (namely, the higher classification accuracy on the forgetting data points) and the lower MIA (Jia et al., 2023) (namely, making it challenging to infer the forgetting identity of a training point), compared to Retrain. Moreover, a higher saliency sparsity yields better testing accuracy. This is not surprising since as shown in (4), the higher saliency sparsity indicates fewer weights of the pre-trained model to be modified during unlearning. Conversely, selecting a lower sparsity ratio may result in over-forgetting due to the excessive number of weights to be modified. In the context of MU for image generation (Fig. A2), a higher sparsity ratio also leads to under-forgetting, while a much lower sparsity ratio causes a rise in FID, introduced by over-forgetting.
SalUn
$10\%$ $30\%$ $50\%$ $70\%$ $90\%$
Forgetting class: “airplane” Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
Non-forgetting classes Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
Figure A3: Image generations of SalUn with different saliency sparsity. The settings follow Fig 4.
Figure A4: The relationship between optimal saliency sparsity and forgetting data amount. The settings follow Table 1.
Additionally, Fig. A3 shows the examples of generated images when forgetting the class ‘airplane’. It’s evident that opting for a 10% saliency sparsity can effectively forget the class ‘airplane,’ but it leads to a decrease in image generation quality. On the other hand, using a 90% saliency sparsity fails to forget the class. In Fig. A4, we present how the ‘optimal’ sparsity level can change vs. the number of forgetting data points, focusing on the case of MU for image classification. The optimal sparsity is determined through a greedy search with a sparsity interval of 10%. It’s important to note that the choice of a 50% sparsity ratio is not universal, and when dealing with a larger amount of forgetting data, a lower sparsity ratio is preferred. This is because a larger forgetting data set requires more weight updates, leading to a lower saliency sparsity.
@@ Appendix C Broader Impacts and Limitations
Figure A5: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color.(More results will be shown in Fig. A6 and Fig. A7)
@@
Figure A6: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color. (Extended results from Fig. A5, more results will be shown in Fig. A7)
@@
Figure A7: Class-wise unlearning results on classifier-free guidance DDPM on CIFAR-10. The forgetting class is marked with a red color. (Extended results from Fig. A6)